A multiple choise test has 3 questions, the first question has 3 possible answers a, b, and c the second question has 4 possible answers a b c and d the 3rd question has 5 possible answers a b c d and e. A student completely does not understand any question and just randomly picks an answer for each question.
A) how many possible choices can he answer with
B) Find the probability that the student answers all 3 questions correctly
C) Find the probability that the studnet answers correctly 1 question and gets the other two questions incorrect.

Answer by stanbon(48568)

the first question has 3 possible answers a, b, and c the second question has 4 possible answers a b c and d the 3rd question has 5 possible answers a b c d and e. A student completely does not understand any question and just randomly picks an answer for each question.
A) how many possible choices can he answer with?: 3*4*5= 60
----------------------
B) Find the probability that the student answers all 3 questions correctly
Prob(3 correct) (1/3)(1/4)(1/5) = 1/60
-----------------------
C) Find the probability that the studnet answers correctly 1 question and gets the other two questions incorrect.
P(RWW or WRW or WWR) = 1/3 + 1/4 + 1/5 = 12/60 = 1/5
==============
Cheers,
Stan H.